# Language-Agnostic Modeling of Source Reliability on Wikipedia

### v2

@@ 3.4.3. User-based features
  • $Proba(R_{add})$, $Proba(R_{rem})$ Probability that, when a domain is added or removed, this is done by the registered users. Contrary to the ratio here we take into account the number of revisions per user.
  • For all user-based features listed above, we compute a version wherein instead of counting any instance of adding a domain, we only count instances of adding it for the first time on an article (“starting”), and similarly counting only the instances that the domain was removed for the last time (“ending”). For example, $Proba(R_{add})$, and $Proba(R_{rem})$ would have a corresponding version $Proba(R_{start})$ and $Proba(R_{end})$ (similarly for the related features).
When we aggregate the various statistics over articles for a domain, we have considered the average and the median and found the best performance with the average, which we use in the models in the following sections. [^6] We have also experimented with versions of the above features that use the aggregation over all URLs in each dataset, instead of domains. However, the resulting features were performing very similarly to those aggregated over domains, and we exclude these from the analysis..
@@ 4. Results
Figure 1. Beeswarm analysis of an XGBoost model trained on Climate Change English dataset (class encoding is 1: reliable and 0: unreliable) and two example classification explanations (score shown is the log-odds ratio).
[…]
[¡short description¿]¡long description¿
@@ 4.1. RQ1: Case study: Climate Change in English
Figure LABEL:fig:RQ2_beeswarm shows the “bee swarm” plot of SHAP (SHapley Additive exPlanations) values that illustrate the importance of each feature in the classifier (Lundberg and Lee, 2017). The top 6 most predictive features (and an aggregation of the rest) are shown. The SHAP values for the top feature $Proba(R_{end})$ – the probability that, when a domain is removed, it is done by a registered editor – reveal that high values of this feature are associated with a negative impact in the domain reliability prediction. In other words, when a registered editor removes a domain, it is more likely to be unreliable. The opposite can be seen for the $\langle SelfPerm_{d}\rangle$ –average proportion of days the domain remains cited since it was first added– and $\langle Perm_{r}\rangle$ –the average number of revisions that the domain remains on the article–: higher values of these features are associated with more reliable sources. $\Sigma CurrPerm_{r}$ –the aggregated number of revisions through which the domain has existed across all articles– also correlates positively with the reliable category, so does as well as the sum of the domain ages across articles, $\Sigma age_{r}$. Finally, “younger” domains – those that have been added to an article more recently, $\langle age_{d}\rangle$ – are more likely to be reliable. This may be a sign of the increased focus of the platform on the quality of sources (Baigutanova et al., 2023a).
@@
In Figures LABEL:fig:RQ1_twitter and LABEL:fig:RQ1_Researchgate we illustrate the performance of the model for two domains of special interest: twitter.com, correctly identified as unreliable, and researchgate.net, incorrectly identified as reliable. In the case of twitter.com, despite being added/present frequently on articles (high $\Sigma CurrPerm_{r}$), it is soon removed by registered editors ($Proba(R_{end})$ and also SelfPermanence, $\langle SelfPerm_{d}\rangle$ is low). In the case of researchgate.net, most of the permanence features point to it being used as a reliable source. This points out that, although this domain is considered unreliable in the perennial source list because the platform does not perform fact-checking or peer reviewing and is considered a self-published source, the Wikipedia community often uses this domain as if it were reliable. In fact, this platform could be used as a way to provide an open-access link to peer-reviewed references when the official version is behind a paywall. Therefore, this specific example shows that our methodology could be used to identify cases where the use of domains is not in line with the established guidelines.
4.2. RQ2: Topic & Language Generalizability
@@
We continue by evaluating the performance of our language-agnostic modeling approach in other topical and linguistic settings. Table 2 shows the F1 macro and class-specific precision and recall for these topics, in the English Wikipedia. We find that, overall, the performance is consistent, with the lowest performance observed for the History dataset, with a recall of 0.69 for the unreliable class. The best performance is achieved on the COVID-19 dataset, possibly due to the ample data on the edits and the more recent nature of the corresponding articles. We find that the amount of edited data is important for the classifier’s performance, as we show later in this section. Finally, we combine all English-language datasets into one, and perform training and LOO validation; the performance of this model is shown in Table 2 under “all topics”. Such a combined model performs on par with other topical models, with especially high precision for the unreliable class (0.89), suggesting that combining knowledge about different topics does not overall degrade the performance of the classifier.
@@
Additionally, we show aggregate statistics of the model performance on the languages in each category in Figure 3. In particular, we are interested in whether our model outperforms the random classifier in terms of F1 macro score (top panel of the figure): we perform this comparison using Mann-Whitney scores and adjust the significance level of 0.05 using the Bonferroni correction for multiple comparisons. For instance, in the Climate Change topic in 100% of the high-resource languages, the models outperform the random baseline, but instead in only 44% of the low-resource languages. The performance on low-resource languages tends to vary widely for different topics. The brown bars show the models for which articles for all topics within a language are used for training and testing. These models perform markedly better than topic-specific ones, echoing our findings for the English language.
Figure 2. Performances (mean $\pm$ stdv) of models trained and validated in languages other than English on the Climate Change datasets. The three batches of languages, separated by a vertical black line, show results on a sample of high (left), mid (centre), and low (right) resource languages.
[…]
[¡short description¿]¡long description¿
Figure 3. Native model performance per topic for different language types. Top panel: % of models that perform better than a random classifier. Bottom panel: Violin plots of corresponding distributions of F1 macro performances.
[…]
[¡short description¿]¡long description¿
Figure 4. Average model F1 macro (line) and standard dev. (shaded area) versus the size of the training dataset in terms of the number of revisions. In English (red), all topics are combined and the data is sampled at regular intervals. The same experiment is repeated, considering only revisions before the implementation of the perennial sources on 2018-07-01 (grey). The performance of models in all other languages (also all-topic) is shown in blue, without sub-sampling.
[…]
[¡short description¿]¡long description¿
@@
The worse performance of our model in the lower-resource languages may be mainly (among others) due to two reasons: the datasets available for training the model are too small, or the behaviour of the users of these languages towards domains is intrinsically difficult to model (their treatment of reliable and unreliable domains is difficult to distinguish). We design two experiments to address these two reasons, starting with the model sensitivity to the amount of training data.
@@
The red line in Figure 4 shows the performance of the model trained on different dataset sizes. In particular, we consider the English (all-topic) dataset and sample the data at log-regular intervals (starting at $10^{3}$, then at $10^{3.25}$ etc. until $10^{6.75}$), resulting in an increasing number of revisions available for training. We repeat this exercise 10 times (mean $\pm$ stdv. shown in the Figure). For comparison, we depict the performance of the models trained on the all-topic datasets in all other languages (shown in blue). Note that in these cases there is no sampling, and for every language, all available data is used. We find that the performance of the English-language model is comparable to that of the other languages when there is little data available (up until about $10^{5}$ revisions) and is about 10% higher than the other languages when almost all data is used for its training. At least in the English-language case, we find that for stable model performance around 10,000 revisions are necessary. At the higher range of data, we find that the classifier’s performance has not levelled off, suggesting that adding more articles (possibly other topics) may improve the performance further.
@@
Turning to editorial behaviour, the introduction of the perennial sources list provides a “natural experiment” to measure a possible editorial behaviour change in response to these additional guidelines. We only examine the impact of this change in the English-language datasets, as it is not clear whether editors from other language editions of Wikipedia used the English perennial sources list. Thus, we created a version of our English-language dataset using only revisions up to 2018-06-30 (i.e., before these guidelines were introduced). To control for the dataset size (which as shown above, may affect model performance), we perform again sub-sampling similar to what we did with the full English-language dataset. The black dash-dotted line and the grey area in Figure 4 show the model performance based on this limited data. We observe its performance to be notably lower than with more recent data (red line and area), indicating that the editorial behaviour before the introduction of the perennial sources list was less predictive of reliable/unreliable domain classes than afterwards. In this case, the performance is similar to that of other languages, showing that that editor behaviour allows extracting meaningful and predictive signals for source reliability even when the editors did not have an official guideline for source reliability. Furthermore, these findings support previous work which shows that the quality of the Wikipedia references improved after the introduction of the perennial sources (Baigutanova et al., 2023a).
4.3. RQ3: Model Adaptation across Topics & Languages
@@
The multilingual nature of Wikipedia presents an opportunity to supplement information available in lower-resource languages with that in the larger ones. Considering our datasets, two directions of adaptation are possible: across languages and across topics. We consider both scenarios in Figure 5, which shows the performance of the models adapted in these two ways: in a cross-language setting (trained on a language and afterwards tested on a different language from the same resourcefulness class, in the same topic) and a cross-topic setting (trained on a topic, tested on a different topic, in the same language). We find that, in both cases, the performance of the adapted models decreases compared to the native performance. The decrease in performance is about the same in cross-language and cross-topic settings. This suggests that the editor’s behaviour around references may be non-identical in different topics, as well as languages. However, again, we see the importance of resourcefulness on the performance: the models adapted across high-resource languages perform better than those across mid and low-resource ones.
@@
Figure 5. Model performance in two adaptation scenarios: cross-language setting (trained on a language, tested on different language from the same class, in the same topic, in red) and a cross-topic setting (trained on a topic, tested on a different topic, in the same language, in green). Results aggregated by language resourcefulness and compared to same-topic and language native models (in blue).
[…]
[¡short description¿]¡long description¿
@@
Finally, we consider all topics together and analyse the performance of our model when trained with different training strategies and tested on each language. In particular, we test whether models trained on higher-resource data would be helpful to those in lower-resource settings. In Figure 6 we aggregate the corresponding results by languages of similar resourcefulness: the top panel shows how for many languages the different models perform significantly better, same as, or worse than a random model (according to Mann-Whitney tests with Bonferroni correction), the center panel shows the same, but compared to the native model, and the bottom panel shows the distribution of F1 macro scores. We find that F1 scores are generally lower when training a model on the English dataset and applying it to low-resource languages, confirming that cross-language adaptation can be problematic, even if a high-resource language such as English is used. However, when training on all high-resource languages or all languages together, the performance for low-resource languages significantly increases. This suggests that including a plurality of languages in training data may capture a variety of user behaviours that are transferable to lower-resource languages.
@@
Nonetheless, each language-specific dataset has its own biases connected to the amount of data available, temporal peculiarities of the articles and their edits, and the volume of editorial activity. To alleviate potential difficulties in adapting feature values between languages, we normalize each dataset using quantile normalization (converting feature values to ranks). This approach further improves the performance of the all-language model adaptation, even on average performing better than the native models. Finally, since the size of the dataset increases with the combination of topics, we train the classifier with the maximum tree depth of 3 instead of 1 but do not find a substantial increase in performance.
@@
Figure 6. Combined Model performances for different training datasets and strategies. Top panel: % of models that perform better, same as, or worse than a random classifier; Center: same, but compared to the native model; Bottom: violin plots of the corresponding distributions of the F1 macro performances.
[…]
[¡short description¿]¡long description¿
4.4. Qualitative Insights on Source Usage
@@
On the other hand, one can use the output of this classifier to find sources that different communities of editors consider inappropriate for their topical domain. For instance, in the Biology dataset, Science-Based Medicine, which is considered generally reliable and has “a credible editorial board” according to the perennial source list, is treated by the editors more similarly to unreliable domains ($P_{r}=0.20$). Similarly, the Huffington Post, rated as “fairly reliable for factual reporting on non-political topics”, scores low in the COVID dataset ($P_{r}=0.07$), possibly due to the politically charged and serious nature of the rhetoric around the pandemic.
4.5. Alternative Sources of Reliability
One limiting factor of this study was the availability of labels for the quality of references. We consider an alternative: Media Bias / Fact Check (MBFC), an “independent website” that provides ratings of “the bias, factual accuracy, and credibility of media sources” [^8] https://mediabiasfactcheck.com/about/, which has been widely used in misinformation research and by other indexes (Barret, 2021; Chowdhury et al., 2020; Ye and Skiena, 2019). We use this credibility metric and consider very high and high as the positive class, and low and very low as negative. With the same features and model as above and LOO cross-validation, we gauge its performance on MBFC. Although the class-average F1 macro is 0.61, the classifier performs especially poorly on the negative class, with a precision of 0.32 and recall of 0.55, suggesting that the features do not capture this metric. Indeed, there may be several reasons why such resources are not suitable for the analysis of Wikipedia references. First, there is little overlap between the sources noted in MBFC and in the English Wikipedia perennial sources list (out of 1156 perennial sources, 272 are in MBFC). Second, the MBFC credibility metric is weakly aligned with the perennial source reliability one. Third, MBFC itself is listed as generally unreliable by the English Wikipedia perennial sources list.

### v3

@@ 3.4.2. Permanence features
  • $\overline{\Sigma CurrPerm_{d}}$, $\overline{\Sigma CurrPerm_{r}}$: Same as above, but considering only the articles where the domain is currently used.
@@ 3.4.3. User-based features
  • $Proba(R_{add})$, $Proba(R_{rem})$ Probability that, when a domain is added or removed, this is done by the registered users. Contrary to the ratio, here we take into account the number of revisions per user.
  • For all user-based features listed above, we compute a version wherein instead of counting any instance of adding a domain, we only count instances of adding it for the first time on an article (“starting”), and similarly, counting only the instances that the domain was removed for the last time (“ending”). For example, $Proba(R_{add})$, and $Proba(R_{rem})$ would have a corresponding version $Proba(R_{start})$ and $Proba(R_{end})$ (similarly for the related features).
We remark that, to aggregate the various statistics over articles for a domain, we have considered both the average and the median and found the best performance with the average, which we use in the models in the following sections. [^6] We have also experimented with versions of the above features that use the aggregation over all URLs in each dataset, instead of domains. However, the resulting features were performing very similarly to those aggregated over domains, and we excluded these from the analysis.
@@ 4. Results
Figure 1. Beeswarm analysis of an XGBoost model trained on Climate Change English dataset (class encoding is 1: reliable and 0: unreliable) and two example classification explanations (score shown is the log-odds ratio). : ¡short description¿
@@ 4.1. RQ1: Case study: Climate Change in English
Figure LABEL:fig:RQ2_beeswarm shows the “bee swarm” plot of SHAP (SHapley Additive exPlanations) values that illustrate the importance of each feature in the classifier (Lundberg and Lee, 2017). The top 6 most predictive features (and an aggregation of the rest) are shown. The SHAP values for the top feature $Proba(R_{end})$ – the probability that, when a domain is removed, it is done by a registered editor – reveal that high values of this feature are associated with a negative impact on the domain reliability prediction. In other words, when a registered editor removes a domain, it is more likely to be unreliable. The opposite can be seen for the $\langle SelfPerm_{d}\rangle$ –average proportion of days the domain remains cited since it was first added– and $\langle Perm_{r}\rangle$ –the average number of revisions that the domain remains on the article–: higher values of these features are associated with more reliable sources. $\Sigma CurrPerm_{r}$ –the aggregated number of revisions through which the domain has existed across all articles– also correlates positively with the reliable category, as does as well the sum of the domain ages across articles, $\Sigma age_{r}$. Finally, “younger” domains – those that have been added to an article more recently, $\langle age_{d}\rangle$ – are more likely to be reliable. This may be a sign of the increased focus of the platform on the quality of sources (Baigutanova et al., 2023a).
@@
In Figures LABEL:fig:RQ1_twitter and LABEL:fig:RQ1_Researchgate we illustrate the performance of the model for two domains of special interest: twitter.com, correctly identified as unreliable, and researchgate.net, incorrectly identified as reliable. In the case of twitter.com, despite being added/present frequently on articles (high $\Sigma CurrPerm_{r}$), it is soon removed by registered editors ($Proba(R_{end})$ and also SelfPermanence, $\langle SelfPerm_{d}\rangle$ is low). For researchgate.net, on the other hand, model and ground truth disagree: most of the permanence features indicate that it is used as a reliable source, while the domain is considered unreliable in the perennial source list because of being a self-publishing platform lacking editorial oversight or peer review. However, as our data points out, Wikipedia editors may still use it pragmatically, for example, to link to freely accessible versions of peer-reviewed papers that are otherwise behind paywalls. This example underscores the potential of our method to reveal discrepancies between Wikipedia’s formal sourcing guidelines and how editors apply them in practice, highlighting edge cases that may warrant further discussion or refinement in policy.
4.2. RQ2: Topic and Language Generalizability
@@
We continue by evaluating the performance of our language-agnostic modeling approach in other topical and linguistic settings. Table 2 shows the F1 macro and class-specific precision and recall for these topics, in the English Wikipedia. We find that, overall, the performance is consistent, with the lowest performance observed for the History dataset, with a recall of 0.69 for the unreliable class. The best performance is achieved on the COVID-19 dataset, possibly due to the ample data on the edits and the more recent nature of the corresponding articles. We find that the amount of edited data is important for the classifier’s performance, as we show later in this section. Finally, we combine all English-language datasets into one and perform training and LOO validation; the performance of this model is shown in Table 2 under “all topics”. Such a combined model performs on par with other topical models, with especially high precision for the unreliable class (0.89), suggesting that combining knowledge about different topics does not overall degrade the performance of the classifier.
@@
Additionally, we show aggregate statistics of the model performance on the languages in each category in Figure 3. In particular, we are interested in whether our model outperforms the random classifier in terms of F1 macro score (top panel of the figure): we perform this comparison using Mann-Whitney scores and adjust the significance level of 0.05 using the Bonferroni correction for multiple comparisons. For instance, in the Climate Change topic, the model outperforms the random baseline in 100% of the high-resource languages, but only in 44% of low-resource languages (in an additional 12% the difference is not statistically significant). The performance on low-resource languages tends to vary widely for different topics. The brown bars show the models for which articles for all topics within a language are used for training and testing. These models perform markedly better than topic-specific ones, echoing our findings for the English language.
Figure 2. Performances (mean $\pm$ stdv) of models trained and validated in languages other than English on the Climate Change datasets. The three batches of languages, separated by a vertical black line, show results on a sample of high (left), mid (center), and low (right) resource languages.: ¡short description¿
Figure 3. Native model performance per topic for different language types. Top panel: % of models that perform better than a random classifier. Bottom panel: Violin plots of corresponding distributions of F1 macro performances.: ¡short description¿
Figure 4. Average model F1 macro (line) and standard dev. (shaded area) versus the size of the training dataset in terms of the number of revisions. In English (red), all topics are combined, and the data is sampled at regular intervals. The same experiment is repeated, considering only revisions before the implementation of the perennial sources on 2018-07-01 (grey). The performance of models in all other languages (also all-topic) is shown in blue, without sub-sampling.: ¡short description¿
@@
The worse performance of our model in the lower-resource languages may be mainly (among others) due to two reasons: the datasets available for training the model are too small, or the behavior of the users of these languages towards domains is intrinsically difficult to model (their treatment of reliable and unreliable domains is difficult to distinguish). We design two experiments to address these two reasons, starting with the model’s sensitivity to the amount of training data.
@@
The red line in Figure 4 shows the performance of the model trained on different dataset sizes. In particular, we consider the English (all-topic) dataset and sample the data at log-regular intervals (starting at $10^{3}$, then at $10^{3.25}$ etc., until $10^{6.75}$), resulting in an increasing number of revisions available for training. We repeat this exercise 10 times (mean $\pm$ stdv. shown in the Figure). For comparison, we depict the performance of the models trained on the all-topic datasets in all other languages (shown in blue). Note that in these cases, there is no sampling, and for every language, all available data is used. We find that the performance of the English-language model is comparable to that of the other languages when there is little data available (up until about $10^{5}$ revisions) and is about 10% higher than the other languages when almost all data is used for its training. At least in the English-language case, we find that for stable model performance around 10,000 revisions are necessary. At the higher range of data, we find that the classifier’s performance has not leveled off, suggesting that adding more articles (possibly other topics) may improve the performance further.
@@
Turning to editing behavior, the introduction of the perennial sources list provides a “natural experiment” to measure a possible editing behavior change in response to these additional guidelines. We only examine the impact of this change in the English-language datasets, as it is not clear whether editors from other language editions of Wikipedia used the English perennial sources list. Thus, we created a version of our English-language dataset using only revisions up to 2018-06-30 (i.e., before these guidelines were introduced). To control for the dataset size (which, as shown above, may affect model performance), we again perform sub-sampling similar to what we did with the full English-language dataset. The black dash-dotted line and the grey area in Figure 4 show the model performance based on this limited data. We observe its performance to be notably lower than with more recent data (red line and area), indicating that the editing behavior before the introduction of the perennial sources list was less predictive of reliable/unreliable domain classes than afterwards. In this case, the performance is similar to that of other languages, showing that editor behavior allows extracting meaningful and predictive signals for source reliability even when the editors did not have an official guideline for source reliability. Furthermore, these findings support previous work, which shows that the quality of the Wikipedia references improved after the introduction of the perennial sources (Baigutanova et al., 2023a).
4.3. RQ3: Model Adaptation across Topics and Languages
@@
The multilingual nature of Wikipedia presents an opportunity to supplement information available in lower-resource languages with that in the larger ones. Considering our datasets, two directions of adaptation are possible: across languages and across topics. We consider both scenarios in Figure 5, which shows the performance of the models adapted in these two ways: in a cross-language setting (trained on a language and afterwards tested on a different language from the same resourcefulness class, in the same topic) and a cross-topic setting (trained on a topic, tested on a different topic, in the same language). We find that, in both cases, the performance of the adapted models decreases compared to the native performance. The decrease in performance is about the same in cross-language and cross-topic settings. This suggests that the editor’s behavior around references may be non-identical in different topics, as well as languages. However, again, we see the importance of resourcefulness on the performance: the models adapted across high-resource languages perform better than those across mid and low-resource ones.
@@
Figure 5. Model performance in two adaptation scenarios: cross-language setting (trained on a language, tested on a different language from the same class, in the same topic, in red) and a cross-topic setting (trained on a topic, tested on a different topic, in the same language, in green). Results aggregated by language resourcefulness and compared to same-topic and language native models (in blue). : ¡short description¿
@@
Finally, we consider all topics together and analyze the performance of our model when trained with different training strategies and tested on each language. In particular, we test whether models trained on higher-resource data would be helpful to those in lower-resource settings. In Figure 6 we aggregate the corresponding results by languages of similar resourcefulness: the top panel shows how for many languages the different models perform significantly better, same as, or worse than a random model (according to Mann-Whitney tests with Bonferroni correction), the center panel shows the same, but compared to the native model, and the bottom panel shows the distribution of F1 macro scores. We find that F1 scores are generally lower when training a model on the English dataset and applying it to low-resource languages, confirming that cross-language adaptation can be problematic, even if a high-resource language such as English is used. However, when training on all high-resource languages or all languages together, the performance for low-resource languages significantly increases. This suggests that including a plurality of languages in training data may capture a variety of user behaviors that are transferable to lower-resource languages.
@@
Nonetheless, each language-specific dataset has its own biases connected to the amount of data available, temporal peculiarities of the articles and their edits, and the volume of editing activity. To alleviate potential difficulties in adapting feature values between languages, we normalize each dataset using quantile normalization (converting feature values to ranks). This approach further improves the performance of the all-language model adaptation, even on average performing better than the native models. In more detail, the model outperforms a random classifier in more than 65% of low-resource settings, and even outperforms native models in 47% of these cases, while being on par with it in 21%. Notably, models trained solely on English fail to generalize well to low-resource languages, performing worse than random in over 70% of cases—highlighting the distinct editorial behaviors across linguistic communities.
Finally, since the size of the dataset increases with the combination of topics, we train the classifier with the maximum tree depth of 3 instead of 1, but do not find a substantial increase in performance.
@@
Figure 6. Combined Model performances for different training datasets and strategies. Top panel: % of models that perform better, same as, or worse than a random classifier; Center: same, but compared to the native model; Bottom: violin plots of the corresponding distributions of the F1 macro performances. : ¡short description¿
4.4. In-Depth Examination of Source Usage
@@
On the other hand, one can use the output of this classifier to find sources that different communities of editors consider inappropriate for their topical domain. For instance, in the Biology dataset, Science-Based Medicine, which is considered generally reliable and has “a credible editorial board” according to the perennial source list, is treated by the editors more similarly to unreliable domains ($P_{r}=0.20$). Similarly, the Huffington Post, rated as “fairly reliable for factual reporting on non-political topics”, scores low in the COVID dataset ($P_{r}=0.07$), possibly due to the politically charged and serious nature of the rhetoric around the pandemic. As editorial policies of sources, as well as perceptions and behaviors of the Wikipedia editors, change, the metrics developed in this study may provide a guide for the improvement and updating of the perennial sources and potential adjustment of the labels in specific domains of expertise.
4.5. Alternative Indicators of Source Reliability
One limiting factor of our study was the availability of labels for the quality of sources. We consider an alternative: Media Bias / Fact Check (MBFC), a third-party website that provides ratings of “the bias, factual accuracy, and credibility of media sources” [^8] https://mediabiasfactcheck.com/about/, which has been widely used in misinformation research and by other indexes (Barret, 2021; Chowdhury et al., 2020; Ye and Skiena, 2019). We use this credibility metric and consider very high and high as the positive class, and low and very low as negative. With the same features and model as above and LOO cross-validation, we gauge its performance on MBFC. Although the class-average F1 macro is 0.61, the classifier performs especially poorly on the negative class, with a precision of 0.32 and recall of 0.55, suggesting that the features do not capture this metric. Indeed, there may be several reasons why such resources are not suitable for the analysis of Wikipedia references. First, there is little overlap between the sources noted in MBFC and in the English Wikipedia perennial sources list (out of 1156 perennial sources, 272 are in MBFC). Second, the MBFC credibility metric is weakly aligned with the perennial source reliability one. Third, MBFC itself is listed as generally unreliable by the English Wikipedia perennial sources list.
In conclusion, MBFC and similar resources are valuable for assessing source reliability in general research contexts. However, the above observations suggest that they are less suitable for evaluating the reliability of sources in the specific context of Wikipedia articles, especially if used as ground truth for supervised classification methods like those presented in this work.